Find three consecutive integers with a sum of 267. Thanks
Answer by checkley77(12844)

LET THE 3 INTEGERS BE X, X+1 & X+2
X+X+1+X+2=267
3X=267-3
3X=264
X=264/3
X=88 IS THE FIRST INTEGER.
88+1=89 IS THE SECOND INTEGER
88+2=90 IS THE LAST INTEGER.
PROOF
88+89+90=267
267=267